# EarthMind: Towards Multi-Granular and Multi-Sensor Earth Observation with Large Multimodal Models

### v1

@@ Appendix F Training Dataset
Table 8: Ablation study on Multi-granular joint training.
@@
Table 9: Ablation study on joint multi-sensor training.
@@
Appendix G More Ablation Studies
Ablation on Multi-Sensor Data Processing. One of the key strengths of EarthMind lies in its ability to handle multi-sensor data beyond standard RGB imagery. To better understand the impact of different preprocessing strategies, we conduct a series of experiments focused on the handling of SAR and multispectral (MS) data. For SAR inputs with fewer than three channels, we compare two strategies: (1) Zero padding, where the missing channels are filled with zeros; (2) Channel replication, where existing channels are duplicated to reach three channels. For multispectral data with more than three bands, we evaluate: (1) Three-band grouping, where every three consecutive bands are grouped to form one RGB-like frame; (2) Single-band grouping, where each band is treated as an individual frame, forming a multi-frame sequence. We evaluate each method under the same training settings and report classification accuracy on the test sets of BigEarthNet (for MS) and SAR Ship Detection (for SAR), as summarized in Tab. 7. Our results show that zero padding outperforms channel replication for SAR data, likely because copying channels introduces redundancy and potential noise. For MS data, single-band grouping slightly improves performance due to better spectral resolution, but incurs substantial computational overhead due to the increased number of tokens. Considering both effectiveness and efficiency, we adopt zero padding for SAR and three-band grouping for MS as our default configuration. In the future, we will consider involve token reduction techniques [71] to reduce overhead cost.
Ablation on Joint Multi-Granular and Multi-Sensor Training. EarthMind is designed to be jointly trained on both multi-granular and multi-sensor data. To validate the effectiveness of this unified training paradigm, we conduct ablation studies along two axes: (1) Multi-Granular Training. We compare two settings: (i) Joint training with image-level, region-level, and pixel-level data. (ii) Independent training for each granularity using the same total amount of data. (2) Multi-Sensor Training. We also compare: (i) Joint training with all available sensor modalities (e.g., RGB, SAR, MS). (ii) Independent training with each modality separately. As shown in Tab. 8 and Tab. 9, the multi-granular co-training strategy consistently outperforms independently trained counterparts, especially on pixel-level tasks. This suggests that high-level semantic supervision (e.g., image-level QA) can improve fine-grained understanding through shared representation learning. Similarly, multi-sensor co-training improves generalization across modalities. Notably, the performance under SAR-only evaluation is significantly enhanced by leveraging complementary information learned from RGB and MS data during joint training. This highlights EarthMind’s ability to exploit cross-modal synergy in both perception and reasoning tasks.
Ablation on Curriculum Training Strategy. EarthMind adopts a curriculum learning strategy that gradually adapts the model from general vision-language data to remote sensing tasks. Specifically, the training proceeds in three stages: (1) pretraining on large-scale natural image VQA and captioning datasets, (2) domain adaptation using EO-specific RGB data, and (3) fine-tuning on multi-sensor data (e.g., SAR). To evaluate the effectiveness of this curriculum, we conduct ablation studies by removing or reordering training stages. As shown in Tab. 10 and Tab. 11, pretraining on general image-language data provides strong foundational capabilities, leading to significant performance improvements on EO tasks, particularly in pixel-level segmentation. Furthermore, RGB-domain training not only enhances performance on RGB inputs but also boosts multi-sensor fusion results, demonstrating its role as an effective bridge between general vision and SAR-specific domains.
Table 10: Ablation study on the first stage of curriculum training.
@@ Appendix G More Ablation Studies
Table 11: Ablation study on the second stage of curriculum learning.
@@
Appendix H More Visualization Results
More visualization results of EarthMind can be seen in Fig. 7.
Figure 7: More visualization of EarthMind.

### v2

@@ Appendix G Training Dataset
Table D: Ablation study on Multi-granular joint training.
@@
Table E: Ablation study on joint multi-sensor training.
@@
Appendix H More Ablation Studies
Ablation on Multi-Sensor Data Processing. One of the key strengths of EarthMind lies in its ability to handle multi-sensor data beyond standard RGB imagery. To better understand the impact of different preprocessing strategies, we conduct a series of experiments focused on the handling of SAR and multispectral (MS) data. For SAR inputs with fewer than three channels, we compare two strategies: (1) Zero padding, where the missing channels are filled with zeros; (2) Channel replication, where existing channels are duplicated to reach three channels. For multispectral data with more than three bands, we evaluate: (1) Three-band grouping, where every three consecutive bands are grouped to form one RGB-like frame; (2) Single-band grouping, where each band is treated as an individual frame, forming a multi-frame sequence. We evaluate each method under the same training settings and report classification accuracy on the test sets of BigEarthNet (for MS) and SAR Ship Detection (for SAR), as summarized in Tab. C. Our results show that zero padding outperforms channel replication for SAR data, likely because copying channels introduces redundancy and potential noise. For MS data, single-band grouping slightly improves performance due to better spectral resolution, but incurs substantial computational overhead due to the increased number of tokens. Considering both effectiveness and efficiency, we adopt zero padding for SAR and three-band grouping for MS as our default configuration. In the future, we will consider involve token reduction techniques (Liu et al., 2025) to reduce overhead cost.
Table F: Ablation study on the first stage of curriculum training.
@@ Appendix H More Ablation Studies
Table G: Ablation study on the second stage of curriculum learning.
@@
Ablation on Joint Multi-Granular and Multi-Sensor Training. EarthMind is designed to be jointly trained on both multi-granular and multi-sensor data. To validate the effectiveness of this unified training paradigm, we conduct ablation studies along two axes: (1) Multi-Granular Training. We compare two settings: (i) Joint training with image-level, region-level, and pixel-level data. (ii) Independent training for each granularity using the same total amount of data. (2) Multi-Sensor Training. We also compare: (i) Joint training with all available sensor modalities (e.g., RGB, SAR, MS). (ii) Independent training with each modality separately. As shown in Tab. D and Tab. E, the multi-granular co-training strategy consistently outperforms independently trained counterparts, especially on pixel-level tasks. This suggests that high-level semantic supervision (e.g., image-level QA) can improve fine-grained understanding through shared representation learning. Similarly, multi-sensor co-training improves generalization across modalities. Notably, the performance under SAR-only evaluation is significantly enhanced by leveraging complementary information learned from RGB and MS data during joint training. This highlights EarthMind’s ability to exploit cross-modal synergy in both perception and reasoning tasks.
Ablation on Curriculum Training Strategy. EarthMind adopts a curriculum learning strategy that gradually adapts the model from general vision-language data to remote sensing tasks. Specifically, the training proceeds in three stages: (1) pretraining on large-scale natural image VQA and captioning datasets, (2) domain adaptation using EO-specific RGB data, and (3) fine-tuning on multi-sensor data (e.g., SAR). To evaluate the effectiveness of this curriculum, we conduct ablation studies by removing or reordering training stages. As shown in Tab. F and Tab. G, pretraining on general image-language data provides strong foundational capabilities, leading to significant performance improvements on EO tasks, particularly in pixel-level segmentation. Furthermore, RGB-domain training not only enhances performance on RGB inputs but also boosts multi-sensor fusion results, demonstrating its role as an effective bridge between general vision and SAR-specific domains.
Appendix I More Visualization Results
More visualization results of EarthMind can be seen in Fig. D.
Figure D: More visualization of EarthMind.